Knowing that each pulley has a radius of 0.3m,
Force \overrightarrow{P} = 29kN
Angle= 50°

using torque method determine:
a) The reaction force \overrightarrow{D} in D.
b) The reaction force \overrightarrow{E} in E
Answer #1

De Ey K EX Рcosp 3.7 8.3 Psing EME=0 Dis 3 Poinß (8-3) - DX (3) =0 Painß (8-3 29 x singso)x8-3 ID 61.462 KN 3 Dx » Efx=0. DxFBDO ACE 4.3 29 kn Ey Ex R 1. Ay 3-7 MA=o 29(4-3) + Ey(6.0) Ey = 15.5875 KN using 15.5875 Dy Dy 22.215 2) 606275 kn Reaction